OpenStudy (shamil98):

Another derivative question.. this time with the quotient rule.. :D x-3/x^2+x+1

OpenStudy (campbell_st):

let \[u(x) = x - 3... so..... u'(x) = 1 \] and \[v(x) = x^2 + x + 1...so.....v'(x) = 2x + 1\] the quotient rule says \[f'(x) = \frac{v \times u' - u \times v'}{v^2}\] just substitute and evalaute

OpenStudy (abb0t):

\(\large\sf \color{red}{\frac{low\frac{d}{dx}(high)-high \frac{d}{dx}(low)}{(low)^2}}\) that helped me learn quotient rule

OpenStudy (abb0t):

low dee high minus high dee low, over low-low. which means: lower function multiplied by derivative of the higher function, MINUS higher function times derivative of the lower function, ALL OVER the lower function twice.

OpenStudy (shamil98):

Hm, soo (x^2 + x + 1)(x-3)' - (x-3)(x^2+x+1)' / (x^2 + x + 1)^2

OpenStudy (shamil98):

(x^2 + x + 1)(1) - (x-3)(2x+1)/(x^2+x+1)^2

OpenStudy (abb0t):

Write it in LaTex form!!! I can barely understand that! But pretty much yes. I think that looks right

OpenStudy (shamil98):

It says to not simplify so that's the final answer.

OpenStudy (shamil98):

well actually (x^2 + x + 1) - (x-3)(2x+1)/(x^2+x+1)^2

OpenStudy (abb0t):

No. That's not the answer. Do not simplify means do not try to cancel out any terms. But you STILL need to find the derivative and perform the algebra on the numerator portion such as distribute and such.

OpenStudy (abb0t):

Because what you're doing above is just tedious algebra, and you SHOULD be very comfortable with algebra if you're doing Calculus now.

OpenStudy (shamil98):

(x^2 + x + 1) - (2x^2 -5x -3) /(x^2+x+1)^2

OpenStudy (shamil98):

WHAT DO YOU MEAN LATEX ? I JUST CHECKED THE SOLUTIONS MANUAL its writtten like this -.- (x^2 + x + 1)(1) - (x-3)(2x+1)/(x^2+x+1)^2

OpenStudy (abb0t):

Use the equation editor button on the bottom!!

OpenStudy (shamil98):

\[\frac{ (x^2 + x + 1) - (2x^2 -5x -3) }{ (x^2+x+1)^2 }\]